# Exploration and Incentivizing Participation in Clinical Trials

### v6

@@ Appendix B Lower bounds for heterogeneous agents
The adversary $\mathtt{adv}$ is constructed as follows. For the first $t_{0}$ rounds, types are drawn from the prior and we use the adversary $\mathtt{adv}_{\psi_{0}}$. For rounds $t>t_{0}$, only type $\theta$ arrives, and we use adversary $\mathtt{adv}^{\dagger}$ from (4.5), determined by the expected sampling probabilities $\left(\,\operatornamewithlimits{\mathbb{E}}[p_{t}]:\,t\in S\,\right)$ (see Lemma C.1). This reduces to the single type estimation problem, and by applying the same argument as in the proof of Theorem 4.5, we have that
@@
The adversary $\mathtt{adv}$ is constructed like in the homogeneous case: $\mathtt{adv}_{\psi_{0}}$ for the first $t_{0}$ rounds, and $\mathtt{adv}^{\dagger}$ from (C.1) afterwards. By the argument from the proof of Theorem 4.5, we have that
@@
Appendix C Statistical estimators
We prove general results for statistical estimators: the lower bound from Eq. (4.5) and the (folklore) upper bound for IPS estimator stated in Eq. (3.3).
C.1 Lower bound: proof of Eq. (4.5)
Let us provide a standalone formulation for Eq. (4.5). Formally, we interpret $\widehat{f}_{\mathtt{mech}}$ as a mapping from arm $a$ and history $\mathtt{hist}_{T+1}$ to $[0,1]$, and call any such mapping an estimator.
Lemma C.1.
@@ Appendix C Statistical estimators
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)\geq\Omega\left(\,\frac{1}{T^{2}}\;\max_{a\in[n]}\;\sum_{t\in S}\min\left\{\,\frac{1}{\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a)\,\right]},\sqrt{|S|}\,\right\}\,\right).$ (C.1)
The adversary in (C.1) only depends on $\widehat{f}_{\mathtt{mech}}$ and the expected sampling probabilities $\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}\,\right]$.
Remark C.2.
@@
Lemma C.1 was known for the special case when $S=[T]$ and each distributions $p_{t}$ is chosen as an independent random draw from some fixed distribution over $\Delta_{n}$ (Dudík et al.,, 2014). The general case, with an arbitrary subset $S$ and a correlated choice of distributions $\left(\,p_{t}:t\in S\,\right)$ may be a tool of independent interest. However, we did not attempt to optimize the constants.
We prove Lemma C.1 in the remainder of this subsection. We use Pinsker’s Lemma to bound the difference in event probabilities.
Lemma C.3 (Pinsker,, 1964).
@@
Letting $P$ be the probability that the absolute difference between the estimator to $\operatornamewithlimits{\mathbb{E}}\left[\,f_{\mathtt{adv}}(a^{*})\,\right]$ is at most $\tfrac{1}{2T}\sum_{t\in S}\delta_{t}$ given adversary $\mathtt{adv}$ and $P^{\prime}$ be the same probability given adversary $\mathtt{adv}^{\prime}$. By Lemma C.3, we have that
$|P-P^{\prime}|\leq\sqrt{\tfrac{1}{2}\;\mathrm{KL}(\mathtt{adv}\|\mathtt{adv}^{\prime})}\leq\textstyle\sqrt{2\;\sum_{t\in S}\;\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]\cdot\delta_{t}^{2}}.$ (C.2)
@@
$\delta_{t}=\left(\,\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]\cdot\sqrt{8\sum_{t\in S}1/\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]}\,\right)^{-1}.$ (C.3)
@@
C.2 Upper bound: proof of Eq. (3.3)
@@
Fixing any oblivious adversary, in the IPS estimator, the random variable $\frac{{\bf 1}_{\left\{\,a_{t}=a\,\right\}}\cdot f(\omega_{a,t})}{p_{t}(a)}$ is independent across different time periods $t$. Since the variance for the sum of independent random variables equals the sum of variance, we have

### v8

@@ Appendix C Lower bounds for heterogeneous agents
The adversary $\mathtt{adv}$ is constructed as follows. For the first $t_{0}$ rounds, types are drawn from the prior and we use the adversary $\mathtt{adv}_{\psi_{0}}$. For rounds $t>t_{0}$, only type $\theta$ arrives, and we use adversary $\mathtt{adv}^{\dagger}$ from (3.5), determined by the expected sampling probabilities $\left(\,\operatornamewithlimits{\mathbb{E}}[p_{t}]:\,t\in S\,\right)$ (see Lemma D.1). This reduces to the single type estimation problem, and by applying the same argument as in the proof of Theorem 3.7, we have that
@@
The adversary $\mathtt{adv}$ is constructed like in the homogeneous case: $\mathtt{adv}_{\psi_{0}}$ for the first $t_{0}$ rounds, and $\mathtt{adv}^{\dagger}$ from (D.1) afterwards. By the argument from the proof of Theorem 3.7, we have that
@@
Appendix D Statistical lower bound: Proof of Eq. (3.5)
We prove the lower bound from Eq. (3.5). Let’s provide a standalone formulation for this result. Formally, we interpret $\widehat{f}_{\mathtt{mech}}$ as a mapping from arm $a$ and history $\mathtt{hist}_{T+1}$ to $[0,1]$, and call any such mapping an estimator.
Lemma D.1.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}\,\right)\geq\Omega\left(\,\frac{1}{T^{2}}\;\max_{a\in[n]}\;\sum_{t\in S}\min\left\{\,\frac{1}{\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a)\,\right]},\sqrt{|S|}\,\right\}\,\right).$ (D.1)
The adversary in (D.1) only depends on $\widehat{f}_{\mathtt{mech}}$ and the expected sampling probabilities $\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}\,\right]$.
Remark D.2.
@@
Lemma D.1 was known for the special case when $S=[T]$ and each distributions $p_{t}$ is chosen as an independent random draw from some fixed distribution over $\Delta_{n}$ (Dudík et al.,, 2014). The general case, with an arbitrary subset $S$ and a correlated choice of distributions $\left(\,p_{t}:t\in S\,\right)$ may be a tool of independent interest. However, we did not attempt to optimize the constants.
We prove Lemma D.1 in the remainder of this subsection. We use Pinsker’s Lemma to bound the difference in event probabilities.
Lemma D.3 (Pinsker,, 1964).
@@
Letting $P$ be the probability that the absolute difference between the estimator to $\operatornamewithlimits{\mathbb{E}}\left[\,f_{\mathtt{adv}}(a^{*})\,\right]$ is at most $\tfrac{1}{2T}\sum_{t\in S}\delta_{t}$ given adversary $\mathtt{adv}$ and $P^{\prime}$ be the same probability given adversary $\mathtt{adv}^{\prime}$. By Lemma D.3, we have that
$|P-P^{\prime}|\leq\sqrt{\tfrac{1}{2}\;\mathrm{KL}(\mathtt{adv}\|\mathtt{adv}^{\prime})}\leq\textstyle\sqrt{2\;\sum_{t\in S}\;\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]\cdot\delta_{t}^{2}}.$ (D.2)
@@
$\delta_{t}=\left(\,\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]\cdot\sqrt{8\sum_{t\in S}1/\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a^{*})\,\right]}\,\right)^{-1}.$ (D.3)
@@
Appendix E Upper bound for IPS: proof of Eq. (2.6)
@@
Fixing any oblivious adversary, since the IPS estimator is unbiased, the mean square error of the IPS estimator equals it variance. Moreover, in the IPS estimator, the random variable $\frac{{\bf 1}_{\left\{\,a_{t}=a\,\right\}}\cdot f(\omega_{a,t})}{p_{t}(a)}$ is independent across different time periods $t$. Since the variance for the sum of independent random variables equals the sum of variance, we have